The Harding Company manufactures skates. The company’s income statement for 20X1 is as follows:

HARDING COMPANY
Income Statement
For the Year Ended December 31, 20X1
Sales (12,500 skates @ $100 each) $ 1,250,000
Variable costs (12,500 skates at $45) 562,500
Fixed costs 400,000
Earnings before interest and taxes (EBIT) $ 287,500
Interest expense 72,500
Earnings before taxes (EBT) $ 215,000
Income tax expense (20%) 43,000
Earnings after taxes (EAT) $ 172,000


a. Compute the degree of operating leverage. (Round your answer to 2 decimal places.)
  



b. Compute the degree of financial leverage. (Round your answer to 2 decimal places.)
Answer #1
Sales 1250000
Variable costs 562500
Contribution margin 687500
a.
Degree of operating leverage = Contribution margin / EBIT = 687500 / 287500 2.39
b.
Degree of financial leverage = EBIT / EBT = 287500 / 215000 1.34
Answer #2

SOLUTION :


a.


Degree of operating leverage (DOL)

= (Sales - Variable costs)/EBIT

= (1250000 - 562500) / 287500

= 566400 / 236400

= 2.3913 

= 2.39  (ANSWER)


b.


Degree of financial leverage (DFL)

= EBIT / EBT 

= 287500 / 215000

= 1.3372

= 1.34 (ANSWER).



answered by: Tulsiram Garg

> Please replace 566400 / 236400 by 687500/287500 in the 4th line of DOL calculation.

Tulsiram Garg Sat, Sep 11, 2021 1:39 AM